A boy is 25 years older than a girl. Five years ago the boy was two times as old as the girl. Find the present age of each.
Answer by macston(5194)

.
G=girl's age; B=boy's age=G+25 yrs
.
B-5=2(G-5)
G+25-5=2G-10
G+20=2G-10
30=G
ANSWER 1: The girl is 30 years old.
.
B+G+25 yrs=30 yrs + 25 yrs=55yrs
ANSWER 2: The boy is 55 years old.
.